INCOME (Dollars) ? вс, 1200 BC2 800 BC, 400 35 40 45 LEISURE (Hours)
INCOME (Dollars) Kate has 80 hours per week to devote to working or to leisure. She is paid an hourly wage and can work at her job as many hours a week as she likes. The following graph illustrates Kate's weekly income-lelsure tradeoff. The three lines labeled BC, BC, and BC illustrate her time allocation budget at three different wages; points A, B, and C show her optimal time allocation choices along each of these constralints BC 1200 BC 800 400
INCOME (Dollars) ? вс, 1200 BC2 800 BC, 400 35 40 45 LEISURE (Hours)
our) For each of the points listed, use the previous graph to complete the following table by indicating the hourly wage at each point and how many hours per week Kate will spend during leisure activities versus working. Wage Leisure Labor Point (Dollars per hour) (Hours) (Hours) A C Based on the data you entered in the previous table, use the orange curve (square symbols) to plot Kate's labor supply curve on the following graph, showing how much labor she supplies each week at each of the three wages 24 22 20 Labor Supply 18 16
WAGE (Dollars per hour) Based on the data you entered in the previous table, use the orange curve (square symbols) to plot Kate's labor supply curve on the following graph, showing how much labor she supplies each week at each of the three wages 24 22 20 Labor Supply 18 16 14 12 10 20 25 30 35 40 45 50 55 60 LABOR (Hours)
WAGE (Dollars per hour) # Homework (Ch 21) 16 14 13 10 4 2 0 510 15 20 35 25 30 LABOR (Hours) 40 45 50 55 60 Suppose that Kate's initial budget line was BC2 and that it then changed to BC3; therefore, Kate's optimal time allocation choice shifted from B to C. As a result of this change, Kate's opportunity cost of leisure ,and she chose to consume leisure. Consequently, in this region, the effect dominates the effect. The corresponding portion of Kate's labor supply curve is
Answer #1
Point Wages (per hour) Leisure(weekly) Work(weekly)
A $ 13.33/hour 35 45
B $ 11.25/hour 40 40
C $ 5.7/hour 45 35

Working notes :-

Wages per hour :-

A = 600/45=13.33

B =450/40=11.25

C =200/35=5.7

Suppose that Kate's initial budget line was BC2 and that has changed to BC3 ; Kate's optimal time allocation choice has shifted from B to C . As a result of change Kate's opportunity cost of leisure has increased as she chose to consume more hours of leisure .Consequently in this region the substitution effect has dominates the consumption effect . The corresponding portion of Kate's labour supply curve is downward sloping.

30+ 12- tu - 10 60 LABOUR Chousu) WA66S CS po hauu